### v1

Abstract
Mutation testing can help minimize the delivery of faulty software. Therefore, it is a recommended practice for developing embedded software in safety-critical cyber-physical systems (CPS). However, state-of-the-art mutation testing techniques for C and C++ software, which are common languages for CPS, depend on symbolic execution. Unfortunately, symbolic execution’s limitations hinder its applicability (e.g., systems with black-box components).
We propose relying on fuzz testing, which has demonstrated its effectiveness for C and C++ software. Fuzz testing tools automatically create test inputs that explore program branches in various ways, exercising statements in different program states, and thus enabling the detection of mutants, which is our objective.
We empirically evaluated our approach using software components from operational satellite systems. Our assessment shows that our approach can detect between 40% and 90% of the mutants not detected by developers’ test suites. Further, we empirically determined that the best results are obtained by integrating the Clang compiler, a memory address sanitizer, and relying on laf-intel instrumentation to collect coverage and guide fuzzing. Our approach detects a significantly higher percentage of live mutants compared to symbolic execution, with an increase of up to 50 percentage points; further, we observed that although the combination of fuzzing and symbolic execution leads to additional mutants being killed, the benefits are minimal (a gain of less than one percentage point).
Keywords:
Mutation testing, Fuzzing, Test data generation
1 Introduction
Ensuring high-quality test suites is essential for quality assurance of embedded software in cyber-physical systems (CPS); indeed, this is the reason why independent software validation and verification activities are required by safety standards ecss40C . Further, although embedded software for CPS can be developed using dedicated visual languages Buffoni2021 , our experience in the space context shows that most software is directly implemented using C and C++, which suggests that approaches supporting validation and verification shall target those languages.
Mutation analysis is an effective method to evaluate the quality of a test suite. It involves measuring the mutation score, which is the proportion of programs with artificially injected faults (mutants) detected by the test suite papadakis2019mutation . There is a strong association between a high mutation score and a high fault-revealing capability for test suites papadakis2018mutation ; Chekam:17 . Additionally, recent studies have demonstrated that mutation analysis can be applied cost-effectively to large CPS software Oscar:TSE .
In practice, mutation analysis warrants the selection of inputs for mutation testing, as test cases should ideally detect all or at least a significant proportion of the generated mutants. A mutant detected by a test suite is considered killed. However, due to the typically high number of mutants generated in large CPS projects, it is challenging for engineers to perform mutation testing manually.
Unfortunately, we lack automated test data generation techniques (automated mutation testing techniques) suitable for the mutation testing of CPS. Indeed, most existing techniques do not target C and C++, which are prevalent in CPS. Further, the state-of-the-art solution for automated mutation testing of C software, SEMu Chekam2021 , is based on the KLEE symbolic execution engine KLEE . While effective for command line utilities, it inherits the limitations of symbolic execution. Specifically, it requires environmental modeling (e.g., network communication) and cannot be applied to programs needing complex analyses for input generation (e.g., programs with floating point instructions). Additionally, it generates test inputs for command line utilities, which are rarely used in CPS, and does not produce unit test cases or target other CPS interfaces.
Search-based techniques developed for other programming languages (e.g., Java fraser2011mutation ) are impractical for C and C++ software due to the difficulty of instrumenting the software to compute dedicated fitness functions (e.g., branch distance). For instance, computing branch distance at runtime necessitates modifying all conditional statements in the software under test (SUT), requiring the source code to be processed with static analysis tools that load all dependencies. Configuring these tools to process the multiple source files in large systems is often impractical unless the tool is well integrated with the compiler used for the SUT. Moreover, CPS source files often rely on architecture-specific C constructs (e.g., for the RTEMS compiler RTEMS ) that are not successfully parsed by static analysis frameworks Oscar:TSE .
Different from search-based testing techniques, grey-box fuzz testing techniques manes2019art (hereafter, fuzzers) can generate test data without relying on complex code instrumentation and thus can be easily applied to C and C++ software. However, fuzzers target console software, while CPS software is usually tested either with system-level test scripts interacting with a hardware emulator or through unit and integration test cases implemented with the same language as the SUT. In this paper, we focus on the automated generation of unit test cases because fuzzing large systems is an open research problem KimXSO22 . Approaches for fuzzing function calls at unit or API level exist FuzzGen ; FUDGE ; APICRAFT ; Jeong2023 ; Hopper , although in the case of large CPS software, their applicability might be limited by static analysis’ scalability issues and difficulties in processing embedded libraries. Further, approaches to generate fuzzing drivers for mutation testing do not exist.
We propose MutatiOn TestIng with Fuzzing (MOTIF), an approach that automatically generates fuzz drivers, leverage the data produced by fuzzers, and exercise a mutated function and the corresponding original function, looking for diverging outputs to detect mutants. MOTIF automatically generates seed files for fuzzers and integrates strategies to process mutant-killing inputs to eliminate false positives due to nondeterminism.
Instead of designing a dedicated fuzzing algorithm, MOTIF leverages state-of-the-art (SOTA) fuzzers because we believe that the fuzzers’ bucket-based coverage strategy can effectively select test inputs that eliminate mutants. This strategy not only selects inputs that lead to different software states but also tracks and is guided by the behavioral differences between the original and mutated functions. Although MOTIF can target any C/C++ program, we designed it taking into account characteristics that are prominent in embedded software for CPS. Specifically, CPS software is not executed from the textual console but directly within the operating system, largely relies on floating point instructions, and leverages specific OS and libraries that complicates its analysis or symbolic interpretation. For these reasons, we present MOTIF as a tool for C/C++ embedded software for CPS.
We introduced MOTIF in a paper presented at the 2023 International Conference on Automated Software Engineering MOTIF where we reported on its effectiveness on three case study subjects in the space domain developed using the C language and demonstrated that it outperforms mutation testing based on symbolic execution. In this paper, we extend our previous work by:
  • Performing a large empirical study to investigate (RQ1) the best fuzzing configurations (compiler, sanitizers, and coverage metric) for mutation testing. Note that the identification of the best fuzzing configuration for different application contexts remains an open problem EURECOM7801 . Our experiments show that the best results are obtained when relying (1) on the Clang compiler, which leads to the quickest fuzz drivers, (2) the address sanitizer (ASAN), which prevents false positives due to violation of function preconditions, and (3) the LAF-Intel LAF:Blog optimization for coverage, which enables computing a better fitness score for mathematical functions, which are, in turn, among the most critical software components. Specifically, the best fuzzing configuration, compared to a standard fuzzing configuration (i.e., AFL++ aflchurn with the GCC GCC compiler), leads to an improvement varying between 1.53 and 4.34 percentage points. Although limited, such improvements may still enable the detection of critical faults; for example, when reusing the generated test cases for regression testing. Further, our findings help engineers make informed decisions on projects setups. For example, mutation testing improvements may not justify a large development endeavor to enable a project to be recompiled with Clang. However, when compiling projects with Clang, ASAN, and LAF-Intel is feasible (e.g., it is sufficient to change a few configuration parameters), the best configuration is a better choice (e.g., enables testing additional boundary cases) and might be considered also for other fuzzing applications beyond mutation testing.
  • Assessing (RQ4) the performance gain achieved by relying on hybrid-fuzzing, which combines fuzzing and symbolic execution. Our results confirm the findings of related work that demonstrate that hybrid fuzzing leads to improved test effectiveness; however, in our experiments, its applicability is limited as it helps significantly improve the mutation score only one in four subjects. The key advantage observed with hybrid-fuzzing is the reduction of killing time (up to 32% faster), which suggests using hybrid-fuzzing for large projects leading to a large number of mutants.
  • Evaluating (RQ5) the cost-effectiveness of reusing mutant-killing inputs to test live mutants; our results suggest that reusing mutant-killing inputs to kill other mutants is recommended when there is a sufficient test budget of at least 5000 seconds per mutant.
  • Replicating, with the best fuzzing configuration, our original MOTIF assessments of the complementarity between fuzzing and symbolic execution for mutation testing (RQ2), and the contribution of MOTIF’s seeding strategy to its results (RQ3). With our subjects, MOTIF kills between 39.5% and 88.7% of mutants, on average, with most of the mutants (52% to 95%) being killed by the search process, not the seed inputs. Further, MOTIF outperforms, by 13.18 to 50.33 percentage points, the results obtained with the mutation testing process based on symbolic execution implemented by the state-of-the-art tool SEMu, thus suggesting that MOTIF is the most appropriate solution for the mutation testing of CPS embedded software. This is the first work comparing fuzzing and symbolic execution for mutation testing.
  • Contributing to the software engineering literature with the first study of mutation testing on industrial software for CPS. In this paper, we assess MOTIF on the control software of ESAILESAILESA , a satellite currently on orbit, and Sentinel-5 UVNS L1b Prototype Processor, a ground software processing radiation data collected by the Sentinel mission of the European Space Agency (ESA) esaSentinel , in addition to the subjects considered in our conference paper: a mathematical library qualified for flight systems, a commercial utility library for nanosatellites, and a serialization/deserialization library, all made available in the context of a project with ESA FAQAS .
  • Providing a technical solution enabling the application of MOTIF to object-oriented C++ programs. Since state-of-the-art mutation testing tools do not target C++ software, our solution enables mutation testing assessment on a broader set of case study subjects.
MOTIF is available online MOTIFGIT ; also, we provide a replication package with our open-source subjects and all our empirical data REPLICABILITY .
The paper proceeds as follows. Section 2 provides background (symbolic execution and fuzzing) and related work (automated mutation testing). Section 3 describes MOTIF. Section 4 presents our extensions to deal with object-oriented programs in C++. Section 5 presents our empirical evaluation. Section 6 concludes the paper.
2 Background and Related Work
This paper relates to the enormous body of work on automated mutation testing and fuzzing; selected, relevant work is discussed below.
2.1 Symbolic execution
Symbolic execution (SE) is a program analysis technique that relies on an interpreter to process the source code of the SUT and automatically generate test inputs Anand2013 . Inputs are represented through symbolic values; during the symbolic execution, the state of the SUT includes the symbolic values of program variables at that execution point, a path constraint on the symbolic values to reach that point, and a program counter. The path constraint is a boolean formula that captures the conditions that the inputs must satisfy to follow that path. Constraint solving SATandCPsurvey:2006 is then used to identify assignments for the symbolic inputs that satisfy the path constraint.
SE presents several limitations, including (1) the need for abstract representations for the external environment and any black-box components used by the SUT—otherwise, the SE engine cannot know what outputs to expect from the environment, (2) path explosion—the SE engine may need to process a large number of paths before satisfying a target predicate, (3) path divergence—abstract representations do not behave like the real systems, (4) the handling of complex constraints, e.g., solving constraints with floating point variables.
A recent solution to partially address the above-mentioned limitations is dynamic symbolic execution (DSE), which consists of treating only a portion of the program state symbolically. Concrete program states help dealing with complex constraints or path explosion (e.g., SE is used after a certain branch has already been reached using a concrete input). However, most frameworks with DSE capabilities, e.g., Angr shoshitaishvili2016state , KLEE KLEE , and S2E S2E , require some degree of environment modeling (e.g., libc library modeling in KLEE), which limits their practical applicability SymCC .
Compilation-based approaches like QSYM yun2018qsym , SYMCC SymCC , SymQEMU poeplau2021symqemu , and SYMSAN SYMSAN augment the original program with instructions to populate and solve symbolic expressions while the original software is executed; such characteristic eliminates some limitations of interpretation-based approaches, thus being applicable to a broader set of software systems. For example, since the symbolic execution interacts with the actual environment there is no need to emulate it within the interpretation layer. SYMCC requires the source code of the SUT, while QSYM relies on dynamic binary instrumentation. SymQEMU, instead, extends the applicability of SYMCC to binary programs by relying on the QEMU emulator for code instrumentation, while SYMSAN relies on dynamic data-flow analysis to reduce the cost of symbolic state management. Since the above-mentioned symbolic execution approaches have shown to provide their best results when combined with fuzzing, a solution referred to as hybrid fuzzing (see Section 2.2), we considered them when assessing the integration of MOTIF with hybrid fuzzing (see Section 5).
2.2 Fuzzing
Fuzzing (or fuzz testing) is an automated testing technique that generates test inputs by repeatedly modifying [^1] To avoid confusion, we avoid the term ‘mutation’ when describing fuzzing techniques. existing inputs; the selection of the inputs to modify is usually driven by metrics collected during the execution of the SUT. Depending on the information collected during program execution, fuzzing techniques (i.e., fuzzers) are classified as black-box, white-box, or gray-box.
In this paper, we focus on grey-box fuzzers because they have demonstrated to effectively maximize code coverage FuzzBenchPaper and discover faults ICST:22:Sarro (mainly crashes and memory errors), two objectives that relate to the problem studied in this paper; indeed, to kill a mutant it is necessary to (1) exercise a mutated statement, which can be achieved by maximizing code coverage, and (2) exercise the mutated statement with many different inputs (i.e., in different states), a common practice in fuzzers to discover crashes and memory errors.
Most fuzzers generate input files to be used for system-level testing of console applications and engineers are therefore required to implement driver programs (hereafter, fuzzing drivers) that rely on the data generated by the fuzzer to test other software interfaces (e.g., APIs, see Section 3.1). Most fuzzers keep a pool of input files and rely on the following evolutionary search process: (1) select an input file from the pool, (2) modify the input file to generate new input files, (3) provide the new input files to the SUT and monitor its execution, (4) report crashes or problems detected through code sanitizers (hereafter, sanitizers UBSan ), (5) add to the pool all the input files that contribute to improve code coverage.
What facilitates the adoption of fuzzers is that they rely on simple dynamic analysis strategies to trace branch coverage of C/C++ programs. A common strategy consists of dynamically identifying branches by applying a hashing function to the identifiers assigned to code blocks by compile-time instrumentation; it is implemented as an extension of popular C/C++ compilers AFL:instr . Further, instead of relying on traditional branch coverage ammann2016introduction , most fuzzers adopt a bucketing approach to track the number of times each branch has been covered by each input file: only once, twice, three times, between four and seven, between 8 and 15, between 16 and 31, etc.; the fuzzers add to the pool those files that cover a bucket not observed before for at least one branch. Such bucketing strategy helps reach software states that are not reachable by simply relying on branch coverage.
Fuzzers mainly differ with respect to the strategy adopted to (1) select what operations to apply in order to modify input files and obtain new ones (e.g., MOpt MOPT relies on a particle swarm optimization algorithm) and (2) select the inputs from the input pool (e.g., AFLfast AFLfast and AFL++ AFL++ rely on a simulated annealing algorithm and prioritize new paths and paths exercised less frequently). Also, fuzzers differ in the strategy adopted to determine interesting inputs. For example, directed grey-box fuzzers WangDGF2020 , instead of maximizing code coverage, aim to reach specific targets, usually a subset of program locations (e.g., modified code) or invalid sequences of operations (e.g., use-after-free).
Hybrid fuzzers Hybrid ; Stephens2016 ; yun2018qsym , instead, rely on grey-box fuzzing to explore most of the execution paths of a program and leverage DSE to explore branches that are guarded by narrow-ranged constraints when the fuzzer does not improve coverage further. Well-known hybrid fuzzing solutions consist of combining AFL AFL with QSYM yun2018qsym and SYMCC SymCC , which have been found to outperform earlier approaches such as Driller Stephens2016 and hybrid testing  Hybrid . The SOTA approach is Fuzzolic FUZZOLIC , which relies on QEMU to generate symbolic queries and solves them with a fuzzing-based technique fuzzySAT .
Some researchers have addressed the problem of generating test drivers to fuzz test program functions as in unit testing FuzzGen ; FUDGE ; APICRAFT ; Jeong2023 ; Hopper ; however, except for Hopper Hopper , they make the assumption that the function under test are already integrated into consumer programs (i.e., programs using the library API) or unit tests Jeong2023 , and all of them rely on complex static and dynamic analyses which are infeasible with large CPS software. Nevertheless, although existing tools do not target the generation of drivers for mutation testing, studying their integration into MOTIF is part of future work.
To avoid relying on program analysis, building on the potential shown by LLM-based test case generation LLMtestingSurvey ; shin2024 ; Fan:2023 , researchers and developers are investigating LLM-based fuzz driver generation Zhang2024 ; OSS:fuzz:llm ; Lyu2024 . For example, a recent study has demonstrated the feasibility of relying on large language models (LLMs) to automatically generate fuzz driver Zhang2024 ; specifically they could generate effective drivers (compile and cover the code) for 91% of the 86 APIs under test with, however, some limitations such as violating API protocol (e.g., parameter initialization) in 39% of the cases. LLM-based generation of mutation testing drivers in MOTIF is part of future investigations.
Other techniques address the problem of generating highly structured input files TensileFuzz ; Skyfire . TensileFuzz generates structured inputs (e.g., image or zip files) by probing random executions to derive constraints for potential input fields, and then relying on string constraint solving to derive inputs TensileFuzz . SkyFire, instead, learns a probabilistic context-sensitive grammar to generate JSON and XML files Skyfire . Such techniques can generate input files with a complex structure but they do not generate unit test cases, which is necessary in our context; however, leveraging those approaches to populate complex data structures may also help with unit-level fuzz testing.
2.3 Automated mutation testing
To kill a mutant, a test case should satisfy three conditions: reachability (i.e., the test case should execute the mutated statement), necessity (i.e., the test case should cause an incorrect intermediate state if it reaches the mutated statement), and sufficiency (i.e., the observable state of the mutated program should differ from that of the original program) offutt1997automatically . Automated mutation testing approaches differ regarding the strategy adopted to satisfy these conditions.
There exist two families of automated mutation testing techniques based respectively on: constraint solving and meta-heuristic search. Only one of them relies on fuzzing brown2020mutation , as further described below.
In this Section, we mainly focus on techniques targeting C and C++ programs because these languages are used in many CPS; unfortunately, the C and C++ languages are more complex to process for static and dynamic analysis techniques than the higher-level languages targeted by most of the techniques in the literature (e.g., Java).
2.3.1 Techniques based on constraint solving
Inspired by the earlier work of Offut et al. offutt1997automatically , Holling et al. execute symbolically the original and mutated functions with input data leading them to generate different outputs holling2016nequivack . A similar technique from Riener et al. riener2011test relies on a bounded model checker (BMC) to select the input values that kill the mutant. Unfortunately, no prototype tools for the above-mentioned approaches are available.
The SOTA tool for automated mutation testing is SEMu Chekam2021 ; SEMUgit , which relies on KLEE to generate test inputs based on SE. To speed up mutation testing, SEMu relies on meta-mutants (i.e., it compiles mutated statements and the original statements together). First, SEMu relies on SE to reach mutated statements (reachability condition). Then, for each mutant, it relies on constraint solving to determine if inputs that weakly kill the mutant exist (necessity condition). For killable mutants, it symbolically runs the mutated and the original program in parallel; when an output statement is reached (e.g., a printf or the return statements of the main function), it relies on constraint solving to identify input values that satisfy the sufficiency condition.
2.3.2 Techniques based on meta-heuristic search
Most of the work on automated mutation testing with meta-heuristic search targets Java software; we report the most relevant techniques below. Ayari et al. ayari2007automatic rely on an Ant Colony Optimization algorithm dorigo2006ant driven by a fitness function that focuses on the reachability condition. Precisely, their fitness measures the distance (number of basic blocks in the program’s control flow graph) between the mutated statement and the closest statement reached by a test case. Fraser and Zeller fraser2011mutation , instead, extended the EvoSuite tool EvoSuite with a fitness function considering the reachability and the necessity conditions (number of statements that are covered a different number of times by the original and the mutated program). The integration of mutation testing into EvoSuite has been further improved with branch distance metrics tailored to the operator used to generate the mutants fraser2015achieving . Recently, EvoSuite has been further extended by Almulla et al. with adaptive fitness function selection (AFFS), a hyperheuristic approach that relies on reinforcement learning (RL) algorithms to determine which composition of fitness functions to use Almulla2022 . Unfortunately, when applied to mutation testing, AFFS does not perform better than SOTA solutions fraser2015achieving .
Concerning C software, we should note the work of Souza et al. souza2016strong , who rely on the Hill Climbing AVM algorithm KorelAVM . They combine three fitness functions that rely on branch distance to measure how far an input is from satisfying each of the three killing conditions. The mutation score obtained with simple C programs ranges between 52% and 93%. The approach has been implemented on top of AUSTIN, a search-based test generation tool for C lakhotia2010austin ; LAKHOTIA2013112 ; AUSTIN ; however, this implementation is not available. A recent search-based testing tool prototype for C is Ocelot Scalabrino:18 ; however, it has not been extended for automated mutation testing. Another key limitation of both Ocelot and AUSTIN is that they implement preprocessing steps that do not work with complex program structures (e.g., we couldn’t apply them to the subject programs considered in our empirical evaluation because of preprocessing errors).
A recent mutation testing technique targeting C software is that of Dang et al. dang2019efficiently , who propose a co-evolutionary algorithm that reduces the search domain at each iteration (the original search domain is replaced by the joint domain of the best solutions found); unfortunately, their prototype is not available.
2.4 Techniques based on fuzzing
The work of Bingham brown2020mutation is the only one to rely on fuzzing to automate mutation testing for C software. For input generation, it relies on TOFU wang2020tofu , a grey-box, grammar-aware fuzzer that generates grammar-valid inputs by modifying existing ones. Similar to Ayari’s work, TOFU’s input generation strategy is guided by the distance between the mutated statement and the closest statement reached by a test case; however, instead of generating unit test cases, it generates input files matching a given grammar. Unfortunately, the results obtained by Bingham are preliminary (they targeted only the Space benchmark SPACE ) and a prototype tool is not available.
DifFuzz DifFuzz , instead, executes two distinct versions of a program and compares their coverage and execution cost (e.g., time) to identify inputs that trigger side channel attacks; although this procedure might be leveraged to perform mutation testing (i.e., execute the original program and the mutant), its implementation targets Java systems. Further it is worth noting that DifFuzz does not compute any difference in terms of code coverage, as well as, the difference in execution time is delegated to a driver that needs to be manually implemented by the software engineer, thus limiting the usability of the approach.
Mu2 MU2 , which has been developed in parallel with MOTIF, is a fuzzer that integrates the findings of search-based unit test generation EvoSuiteMutation to generate test input files with fuzzing: it relies on the mutation score to drive the generation of test inputs. Different from MOTIF, which tests each mutant independently from the others, potentially with different inputs, Mu2 tests every live mutant with each generated input and, in the input pool, prioritizes those inputs that increase the mutation score. Results show that Mu2 kills more mutants than the inputs generated by a traditional fuzzer to test the original function. However, it is unclear whether Mu2’s approach (i.e., testing all the live mutants together) is more effective than that of MOTIF (i.e., executing a fuzzer to test the original and mutated function in sequence). Unfortunately, a direct comparison of Mu2 and MOTIF is not feasible because the scalability of Mu2 is enabled by dynamic classloading and instrumentation, two options that are feasible for Java programs but not for the C/C++ programs targeted by MOTIF. Further, by targeting Java, Mu2 can easily determine if mutants are killed by relying on the method ‘equals’, which is implemented by every class to determine if two instances are equal; the method ‘equals’ is not available in C and C++ software. Mu2’s results follow previous work showing that, in Java benchmarks, prioritizing inputs that increase the mutation score may lead to higher branch coverage and mutation score than traditional prioritization strategies based on branch coverage CoverageGuided .
2.5 Industrial applications of mutation testing
Most mutation testing approaches target Java software; the few mutation testing approaches targeting C/C++ software have been assessed with open source console utilities Chekam2021 or simple algorithm implementations (e.g., merge sort) holling2016nequivack ; riener2011test , but there is no work assessing mutation testing tools on C/C++ industrial software. Last, most empirical studies on large industrial software systems concern mutation analysis Just2021Practical , not mutation testing.
2.6 Summary
To summarize, our research is motivated by the lack of support for automated mutation testing of C/C++ software. The SOTA approach for the automated mutation testing of C/C++ software (i.e., SEMu) relies on KLEE and inherits its limitations, making it inapplicable to most CPS software; further, it does not generate unit test cases but selects inputs for console programs. Other SE tools (QSYM and SYMCC) also present technical limitations preventing their application to CPS software. Search-based approaches for the mutation testing of C/C++ software present acute feasibility challenges due to static analysis, which is needed for branch distance fitness but does not scale in large software projects. Though fuzzing appears to be a feasible input generation strategy for mutation testing, existing fuzzers do not generate test drivers for unit testing. The only fuzzer proposed for mutation testing is not available for download and its results are very preliminary.
Last, our paper addresses the lack of empirical assessments of mutation testing on industrial software by targeting libraries and control software used in satellite systems as case studies.
3 Proposed approach: MOTIF
Figure 1: The MOTIF process.
Inspired by the work of Holling et al. holling2016nequivack , MOTIF aims to identify a set of test inputs that lead to different outputs when given to both the original and mutated functions. To achieve such objective with fuzzing, MOTIF generates a fuzzing driver for each mutated function. The fuzzing driver reads the input data generated by the fuzzer and supplies it as arguments to both the original and mutated functions. Finally, the fuzzing driver compares the outputs from both functions. If the outputs differ, the mutant is considered killed.
Our intuition is that fuzzers might be effective at killing mutants because, by invoking both the original and the mutated functions within the same fuzzing driver, we can leverage the bucket-based fuzzing strategy to cover the different branches in the two functions in diverse ways, thus reaching those program states that enable killing mutants. Essentially, the bucket-based fuzzing strategy may help kill mutants by preserving, during test generation, inputs that lead to incorrect intermediate states but do not kill the mutant (i.e., they do not meet the sufficiency condition). Subsequent iterations of the fuzzer’s evolutionary search process (see Section 2.2) may modify these inputs so that they not only reach an incorrect intermediate state but also meet the sufficiency condition. Indeed, if differences in coverage between the original and mutated functions are observed, it indicates that the functions behave differently, resulting in different outputs and the mutant being killed schuler2013covering ; schuler2009efficient ; Oscar:TSE . Additionally, significant differences in coverage lead to new buckets being covered, and since fuzzing favors inputs that cover new buckets, it indirectly leads to inputs that kill mutants. We leave to future work the extension of fuzzers with dedicated strategies; for example, instead of measuring the coverage of the mutated function, the fuzzer could measure the difference in coverage between the original and the mutated function, and use this information to prioritize the inputs in the fuzzer queue (e.g., testing inputs that leads to larger differences first).
MOTIF creates all the necessary scaffolding to test both the original and mutated functions, and to compare their outputs. Specifically, MOTIF follows the workflow illustrated in Figure 1, which comprises the five steps detailed below.
MOTIF takes as input a set of mutants (source files) to be killed; each mutant matches the original source file except for the statements modified by a mutation operator. The MOTIF Steps in Figure 1 are repeated for each mutant. However, Steps 1 and 2 can be executed only once for all the mutants belonging to the same function; indeed, the structure of the input and output data of a function is not changed by mutation—we do not target interface mutation delamaro2001interface .
3.1 Step 1 – Generate the fuzzing driver
In Step 1, MOTIF relies on the clang static analysis library CLANG to analyse the SUT and determine the types of the parameters required by the function under test. This information is then used to generate a fuzzing driver for mutation testing; an example fuzzing driver for the function T_POS_IsConstraintValid belonging to our ASN1lib case study subject is shown in Listing 1. The fuzzing driver renames the mutated function by adding the prefix mut_.
1int main(int argc, char** argv){
2 load_file(argv[1]); // load the input file and
3 // extends the input with random data if needed
4
5 /* Variables for the original function */
6 T_POS origin_pVal; // for the first parameter
7 int origin_pErrCode; // for the second parameter
8 /* Variables for the mutated function */
9 T_POS mut_pVal; // for the first parameter
10 int mut_pErrCode; // for the second parameter
11 /* Variables for the return values */
12 flag origin_return; // for the original
13 flag mut_return; // for the mutant
14
15 /* Copy the input data to the variables for the original function */
16 get_value(&origin_pVal, sizeof(origin_pVal), 0);
17 get_value(&origin_pErrCode,sizeof(origin_pErrCode),0);
18 log("Calling the original function");
19 origin_return = T_POS_IsConstraintValid(&origin_pVal, &origin_pErrCode);
20
21 /* Copy the same input data to the variables for the mutated function */
22 seek_data_index(0); //reset the input data pointer
23 get_value(&mut_pVal, sizeof(mut_pVal), 0);
24 get_value(&mut_pErrCode, sizeof(mut_pErrCode), 0);
25 log("Calling the mutated function");
26 mut_return = mut_T_POS_IsConstraintValid(&mut_pVal, &mut_pErrCode);
27
28 log("Comparing result values: ");
29 ret += compare_value(&origin_pVal, &mut_pVal, sizeof(origin_pVal));
30 ret += compare_value(&origin_pErrCode,&mut_pErrCode, sizeof(origin_pErrCode));
31 ret += compare_value(&origin_return, &mut_return, sizeof(origin_return));
32
33 if (ret != 0){
34 log("Mutant killed");
35 safe_abort();
36 }
37 log("Mutant alive");
38 return 0;
39}
Listing 1: Example fuzzing driver for the ASN1lib subject.
The fuzzing driver contains two sets of variables (Lines 5-7 and 8-10) whose types match the parameters of the function under test and are provided as input to both the original and the mutated function. In our example, it declares a struct T_POS and an int variable. These two sets of variables are then assigned by performing a byte-by-byte copy of a same portion of input file provided by the fuzzer (Lines 16-17 and 23-24, achieved by the function get_value); MOTIF ensures to copy a number of bytes to match the size of the assigned variable. If the input file provided by the fuzzer is shorter than required (the file modifications performed by fuzzers include shortening files), MOTIF extends it with random data (Line 2). Additionally, the fuzzing driver declares the variables required to store the functions’ return values (Lines 11 to 13).
The original and the mutated functions are then invoked (Lines 19 and 26). The fuzzing driver then compares the outputs generated by both functions (Lines 28-31). Unfortunately, in C and C++, distinguishing between input and output parameters is complicated by the presence of pointer and reference arguments. Additionally, determining input parameters through data-flow analysis is impractical, as it requires preprocessing the SUT with a static analysis framework (e.g., LLVM LLVM ), which is often not feasible for CPS software Oscar:TSE . Therefore, we use a straightforward approach to compare outputs which consists of comparing all parameters and return values of the original and mutated functions; such approach does not lead to incorrect mutant killing because input parameters remain unmodified. For pointers, we compare the data they point to (e.g., an int instance for int*). If the pointer is used as an array, the end-user can specify the expected length of the array, so the array data can be compared. When arrays are inputs to the function under test, the end-user may not need to provide the length, as MOTIF automatically generates arrays with a default length of 100. If the function under test dynamically allocates arrays, the end-user should specify the minimal possible length (e.g., an array of length one) to avoid false positives from out-of-bounds readings. For data structures with pointer fields, the pointed data length and initialization procedure can be specified. If the outputs differ, the fuzzing driver halts execution with an abort signal (Line 35 in Listing 1), allowing the fuzzer to detect the aborted execution and store the input file. MOTIF then stops the fuzzer because the mutant has been killed.
Table 1: Seeds assigned to types
Type Seed 1 Seed 2 Seed 3
int -1 0 1
Bool False True
float -3230283776.0 0.0 1072693248.0
double 13826050856027422720.0 0.0 4602891378046628864.0
char 0xFF 0x00 0x41
byte 0xFF 0x00 0x41
ISO8601 2145916800.999999999 1970-01-01T00:00:00Z 2038-01-01T00:00:00Z
3.2 Step 2 – Generate seed files
In Step 2, MOTIF creates seed files based on the input parameter types for the function under test. These seed files are used by the fuzzer to initiate the testing process. Typically, fuzzers are executed with seed files that correspond to typical inputs for the SUT. In our approach, we automatically generate seed files that contain enough bytes to populate all input parameters with values that cover basic cases. Specifically, for each primitive type, we have identified three representative seed values for typical input partitions, as shown in Table 1. For instance, for numeric values, we provide zero, a negative number, and a positive number. Using these seed values, MOTIF generates up to three seed files for each fuzzing driver, ensuring that each seed value is covered at least once for every input parameter.
Example seed files for function T_POS_IsConstraintValid are provided in Figure 2 (type definitions in Listing 2). MOTIF can also generate seed files for complex input types. For instance, the struct T_POS received as input by function T_POS_IsConstraintValid consists of an enum (named kind), which specifies the type of data stored inside the rest of the struct, and a union (named u), which is sufficiently large to contain the data for all the data types selectable with the variable kind. MOTIF treats such struct as an int array thus filling it with the seeds 0xFFFFFFFF, 0x00000000, and 0x00000001. The first four bytes in the seed files (see Figure 2) belong to the enum item kind, and are filled with the seed values of the int type. The same happens for the union field u but, since the union has a size of 8,052 bytes (size of subTypeArray with 4 bytes padding [^2] https://research.nccgroup.com/2019/10/30/padding-the-struct-how-a-compiler-optimization-can-disclose-stack-memory/), MOTIF repeats the same set of four bytes 2,013 times. The last four bytes belong to the second parameter of T_POS_IsConstraintValid, the int *_pErrCode.
Figure 2: Seed files generated for the fuzzing driver in Listing 1.
1typedef enum { T_POS_NONE, longitude_PRESENT,
2 latitude_PRESENT, height_PRESENT,
3 subTypeArray_PRESENT, label_PRESENT,
4 intArray_PRESENT, myIntSet_PRESENT,
5 myIntSetOf_PRESENT, anInt_PRESENT
6} T_POS_selection;
7
8typedef struct {
9 T_POS_selection kind;
10 union { asn1Real longitude; asn1Real latitude;
11 asn1Real height; My2ndInt anInt;
12 T_POS_label label; T_ARR intArray;
13 T_SET myIntSet; T_SETOF myIntSetOf;
14 T_POS_subTypeArray subTypeArray;
15 } u;
16} T_POS;
Listing 2: Definition of struct T_POS
3.3 Step 3 – Compile the SUT
In Step 3, MOTIF produces an executable fuzzing driver by compiling the source code of the fuzzing driver, the mutated function, and the SUT using the fuzzer compiler, which is essential for gathering the code coverage information needed by the fuzzer. This step also compiles a false positive driver (see Section 3.4 below).
3.4 Step 4 – Perform mutation testing
In Step 4, MOTIF runs the fuzzer to generate inputs for the executable fuzzing driver; in our experiments we leverage the AFL++ fuzzer but the approach is generic and can work with any grey-box fuzzer for C/C++. The fuzzer keeps generating input files until it reports one or more crashes, after which MOTIF halts the fuzzer. This process leads to the generation of fuzzing driver logs and crashing inputs (i.e., input files that caused a crash during the execution of the fuzzing driver). Since fuzzers generate several input files from each input taken from the file pool, and since all of them are executed by the fuzzer, more than one crashing input may be reported.
Fuzzing driver logs include checkpoints indicating the progress of testing (see Lines 18, 25, 28, 34, 37 in Listing 1). For each crashing input, MOTIF processes the corresponding logs to distinguish between:
  • Crashes occurring during the execution of the original function. They indicate either the presence of a fault in the original function or the violation of preconditions. We ignore these inputs because they do not correspond to inputs killing a mutant.
  • Crashes occurring during the execution of the mutated function. Since the mutated function is executed after the original one, we can safely conclude that the test inputs do not cause any crash in the original function. Therefore, the observed crashes indicate that the mutant introduced a fault that was exercised by the input. Thus, we can conclude such inputs kill the mutant.
  • Aborted executions due to the fuzzing driver determining that the mutant has been killed (see Line 35 in Listing 1).
MOTIF retains all test inputs that kill a mutant. However, the function under test may produce non-deterministic outputs meaning that, despite observed differences in outputs, the inputs may not have actually killed the mutant. For instance, two consecutive invocations of a function that reads and writes global variables may lead to different outputs even if the mutated statement is not exercised; consequently, the input suggested by the fuzzer would be a false positive. To minimize false positives, MOTIF automatically re-executes every test input killing a mutant using a modified version of the fuzzing driver that invokes the original function instead of the mutated function. If this false positive driver, as we refer to it, reports a difference in outputs of the two function calls, it implies that the function under test is non-deterministic and thus that the input may not kill the mutant. MOTIF considers mutants exclusively killed by false positive inputs to be live. To kill mutants in functions that modify global state variables, the user must manually add, in the fuzzing driver between the two function calls, instructions to reset each state variable, similar to other fuzzing approaches for unit and library testing (e.g., LibFuzzer LibFuzzer ).
3.5 Step 5 – Test the SUT
In this Step, MOTIF generates a test driver for the SUT. An example test driver for function T_POS_IsConstraintValid is shown in Listing 3. The test driver is identical to the fuzzing driver except that (1) it only invokes the original function (Line 5 in Listing 3), (2) instead of comparing outputs from two function invocations, it prints the output data generated by the original function (Lines 7 to 9), and (3) it includes assertions that compare the execution results of the original function with the expected results (expected results are the ones observed during fuzzing for the original function). The test driver is used to test the original SUT with the inputs that kill the mutant and the outputs should then be verified by a software engineer based on the SUT specifications. If the observed output values are correct, they can be used as oracles for future regression testing, which is the reason why the test driver already includes automatically generated test assertions (i.e., MOTIF generates a complete regression test case). Otherwise, a fault has been identified in the SUT. This scenario highlights one of the key advantages of mutation testing: When testing the SUT with inputs that detect simulated human mistakes (mutants), actual faults in the SUT are more likely to be discovered than with randomly selected inputs (e.g., in our experiments we discovered five faults in our subjects).
In our test driver, the generation of print statements for structs and pointers is determined by the configuration of the fuzzing drivers. By default, all bytes within a struct are printed. When pointers are involved, if the end-user specifies the size of the data to which the pointers refer, the test driver prints the data pointed to, rather than the pointer value.
1int main(int argc, char** argv){
2 load_file(argv[1]); /* load the input file */
3 // Declaration of variables and assignment with input file data missing to save space...
4 /* Invoke the original function*/
5 _return = T_POS_IsConstraintValid(&pVal, &pErrCode);
6 /* Print output values of the original function */
7 printf_struct("pVal (T_POS)=", &pVal, sizeof(pVal));
8 printf("pErrCode (int) = %
9 printf("return (flag) = %
10 return 0;
11}
Listing 3: Example test driver for the ASN1lib subject.
4 MOTIF C++ extensions
In this Section, we describe how we extended MOTIF to test C++ programs. This extension is necessary to accommodate the four key object-oriented programming (OOP) properties – abstraction, encapsulation, inheritance, and polymorphism – and their impact on mutation testing with MOTIF. Note that these problems are specific to C++ because mutation testing approaches for higher-level programming languages such as Java MU2 ; fraser2011mutation can leverage reflection to overcome them, as described below. For the sake of clarity, in this Section, we call MOTIF-C the version of MOTIF that handles C code and was presented in Section 3, and MOTIF-C++ the version of MOTIF extended to deal with C++ features. Listing 4 provides a running example.
Abstraction is a key property of grouping variables and operations that are highly related. Classes are the means to declare those variables (i.e., attributes) and operations (i.e., methods) as new data types. It reduces the complexity of developing software by letting users use the classes without knowing how its features are implemented. Instead, for testing, instantiating a class is necessary to exercise the class instance methods. But MOTIF-C does not instantiate classes since, in C, operations are implemented by functions, which do not require any object to be instantiated. In C++, static methods play the same role as C functions, however they are rare. Consequently, MOTIF-C cannot be used to test most C++ methods.
1class SatelliteModule {
2 private:
3 /** Charge in watts per hour **/
4 double getCharge(){ return power; }
5
6 protected:
7 /** Average consumption in watts per hour **/
8 double getAverageConsumption(){
9 return consumedWatts()/elapsedTime();
10 }
11
12 public:
13 /** Hours before being out of charge **/
14 double getAutonomy(){
15 return getCharge()/getAverageConsumption();
16 }
17};
18
19class ChargeableModule : public SatelliteModule {
20 private:
21 /** Average watts charged per hour **/
22 double getAverageCharging(){
23 return chargedWatts()/elapsedTime();
24 }
25
26 protected:
27 double getAverageConsumption(){
28 return super::getAverageConsumption()-getAverageCharging();
29 }
30
31 public:
32 ChargeableModule(double consumptionBoundary, double chargeBoundary){
33 this.consumptionBoundary = consumptionBoundary;
34 this.chargeBoundary = chargeBoundary;
35 }
36
37 /** True if the mission lasts less than the autonomy of the module**/
38 bool isAutonomous( Mission mission){
39 return mission.duration() <= getAutonomy();
40 }
41};
42
43/** Absract class for missions **/
44class Mission {
45 public:
46 /** Duration in hours **/
47 virtual double duration();
48}
49
50class StaticMission {
51 private:
52 double presetDuration;
53
54 public:
55 StaticMission(double presetDuration){
56 this->presetDuration=presetDuration;
57 }
58
59 double duration(){
60 return presetDuration-elapsedTime();
61 }
62}
63
64template <class ObjectiveT> class DynamicMission {
65 public:
66 void objectiveAchieved(ObjectiveT T){
67 //remove from the list of objectives to achieve
68 this->objectives.remove(T);
69 }
70
71 double duration(){
72 return objectivesToAchieve()/objectivesAchievedPerHour();
73 }
74}
Listing 4: C++ example classes. For simplicity  we show only in-class method declarations and we hide some method declarations.
Encapsulation is realized by hiding (i.e., making private or protected) a subset of the instance variables and methods implemented by a class. Consequently, a subset of class methods (i.e., private and protected methods) cannot be exercised by MOTIF-C; indeed, functions that do not belong to a class (e.g., the main function of the MOTIF fuzzing driver), cannot access the class private and protected members.
Inheritance enables a class to reuse (i.e., inherit) the methods declared by its superclass, if any. Inheritance complicates the unit testing automated by MOTIF-C because it leads to dependencies between superclasses and subclasses. Indeed, not only a subclass may execute superclass methods but the execution of a method declared by a superclass may trigger the execution of a method in a subclass. For example, the execution of method ChargeableModule.isAutonomous() in Listing 4 triggers the execution of the superclass method SatelliteModule.getAutonomy(), which, in turn, triggers the execution of ChargeableModule.getAverageConsumption() and ChargeableModule.getAverageCharging(). As a consequence, the solution adopted by MOTIF-C to test mutants (i.e., creating a copy of the original method and renaming it) would prevent the testing of methods following such pattern; for example, renaming a mutant of method getAverageConsumption() in ChargeableModule as mut_getAverageConsumption() would prevent its call from method SatelliteModule.getAutonomy(), thus rendering testing ineffective.
Polymorphism refers to multiple language features. Ad hoc polymorphism (i.e., method overloading) indicates that methods with a same name can process arguments of different types. Since these methods have distinct implementations, they do not affect MOTIF-C because it treats them as distinct functions. Parametric polymorphism (i.e., class templates) implies that certain methods can be exercised only if the object they belong to has been instantiated with appropriate parameters (i.e., a specific class name). This is the case for class DynamicMission, which requires the specification of the data type used to model mission objectives as in “mission = new DynamicMission<Landmark>();Subtype polymorphism occurs when the code contains a call to a method belonging to a certain object, but the concrete method to be invoked is decided at runtime, depending on the object type. This is the case for method ChargeableModule.isAutonomous(), which invokes Mission.duration() and, at runtime, may execute the implementation of duration() provided by either StaticMission or DynamicMission. To test those methods, it is necessary to determine the type of the object to be instantiated and passed as argument, which is not supported by MOTIF-C.
To address the issues above, MOTIF-C++ implements a set of extensions that are summarized in Table 2 and described below.
Table 2: MOTIF-C++ extensions to deal with C++ language features.
Language feature MOTIF-C++ extension
Inheritance Copy mutant class.
Abstraction Automatically generate test utility class.
Ad hoc polymorphism Separate testing for each method.
Encapsulation Manual editing of test utility class.
Parametric polymorphism
Subtype polymorphism
1class mut_ChargeableModule : public SatelliteModule {
2
3 //... same code as in ChargeableModule
4
5 bool isAutonomous( Mission mission){
6 //MUTANT 1: replaced <= with <
7 return
8 mission.duration() < getAutonomy();
9 }
10};
Listing 5: Mutant class for a mutant of method isAutonomous.
1class mut_ChargeableModule : public SatelliteModule {
2
3 //... same code as in ChargeableModule
4
5 protected:
6 double getAverageConsumption(){
7 //MUTANT 2: removed ’- getAverageCharging()’
8 return super::getAverageConsumption();
9 }
10
11 //... same code as in ChargeableModule
12};
Listing 6: Mutant class for a mutant of method getAverageConsumption.
To deal with inheritance, MOTIF-C++ injects the mutated method into a copy of the original class. Examples are shown in Listing 5 and 6, which present two distinct mutants of class ChargeableModule. Note that in the case of Listing 6, the creation of a renamed class copy, instead of a renamed method copy as in MOTIF-C, ensures that the mutated method is executed when testing mut_ChargeableModule.isAutonomous(). In Java, this issue is addressed in a similar but more efficient way MU2 , by leveraging the classloader mechanism to load in memory both the hierarchy of the original class and its mutated version (one for each mutant), which is not feasible in C++ because of the lack of a classloader managed by a virtual machine.
To deal with abstraction, MOTIF-C++ creates a fuzz driver that, instead of invoking the mutated method, creates an instance of a test utility class to which it delegates the instantiation of the class under test and the execution of the method under test. An example fuzz driver is shown in Listing 7; except for the instantiation of the utility class, it matches the fuzz driver created by MOTIF-C. A test utility class (i.e., MOTIF::TestUtil) is shown in Listing 8. Test utility classes are automatically generated by MOTIF-C++ in Step 1, but often need to be manually modified as described in Section 4.1. A utility class includes state variables to hold the values to be passed as arguments to the constructor of the class under test and to the method under test (Lines 4-9 in Listing 8). Still in MOTIF::TestUtil, two additional state variables are used to store references to the value returned by the method under test (Line 11) and the instance under test (Line 13). The solutions implemented by approaches targeting Java never need manual intervention because they can leverage reflection to determine what are the constructors and, recursively the constructors for their parameters, and just rely on automated trial and error to identify appropriate initializations MU2 ; fraser2011mutation .
In MOTIF::TestUtil, methods call and compare take care of the execution of the method under test and the comparison of the results obtained by the original and the mutated method, respectively. Method call (Lines 13-25, Listing 8), which executes the method under test, includes automatically generated instructions to populate parameters with fuzzed data (Lines 17 and 18). If the constructor of a function under test receives as input an abstract type, the instructions needed to construct the concrete type to be used shall be manually added (Lines 22-24). Last, method call invokes the method under test and stores the returned value (Line 26). Method compare, instead, is invoked by the fuzzing driver (Line 13, Listing 7) to compare the results obtained by the original and the mutated class after the execution of the mutant; its logic follows the one of the fuzzing driver generated by MOTIF-C. Approaches for Java MU2 ; fraser2011mutation can leverage programming conventions and reflection to compare object states by leveraging the equals method, which is normally overridden to compare two instances of a class, or inspector methods, which can be identified through naming conventions or purity analysis (simpler to achieve on Java bytecode Salcianu2005 ).
To deal with protected and private methods (i.e., encapsulation), instead of modifying the class under test to specify the test utility class as a friend class, we suggest the end-user to exercise, in the fuzzing driver, a public method that invokes the private/protected method under test. For example, to test mutant 2 in Listing 6, the end-user may still rely on the fuzzing driver shown in Listing 7, because it tests method isAutonomous, which, in turn, executes method getAverageConsumption (i.e., the mutated method). In Java, protected and private methods can be invoked through reflection Arcuri2017 .
1int main(int argc, char** argv){
2 load_file(argv[1]); // load the input file and
3 // extends the input with random data if needed
4
5 log("Calling the original function")
6 auto origin = new MOTIF::TestUtil<ChargeableModule>();
7 origin->call();
8
9 log("Calling the mutated function")
10 auto mutant = new MOTIF::TestUtil<mut_ChargeableModule>();
11 mutant->call();
12
13 if (orig->compare(mut) != 0){
14 log("Mutant killed");
15 safe_abort();
16 }
17 log("Mutant alive")
18}
Listing 7: Example MOTIF-C++ fuzz driver.
1namespace MOTIF{
2template<typename TypeA>
3class TestUtil{
4 //parameters for the constructor
5 double constructor_1;
6 double constructor_2;
7 //parameters for the method under test
8 Mission parameter_1;
9 int manual_1;
10 //return value
11 bool _return;
12 //instance
13 TypeA object;
14
15 void call(){
16 //populate variables with fuzz data
17 get_value(&constructor_1, sizeof(constructor_1) );
18 get_value(&constructor_2, sizeof(constructor_2) );
19
20 //instantiate the class under test
21 this->object = new TypeA(param1, param2);
22 //instructions for parameters of an abstract type
23 get_value(&manual_1,sizeof(manual_1));//added
24 this->parameter_1 = new StaticMission(manual_1);
25 //invoke the method under test
26 this->_return = object->isAutonomous(this->parameter_1);
27 }
28
29 template<typename TypeB>
30 int compare(TestUtil<TypeB> *rhs){
31 int ret = 0;
32 ret += compare_value(this->parameter_1,rhs->parameter_1);
33 ret += compare_value(this->_return,this->_return);
34 return ret;
35 }
36}
37}
Listing 8: Example utility class.
4.1 Manual tuning of MOTIF’s test utility class
Like MOTIF-C, MOTIF-C++ automatically identifies the variables to create in the test utility class based on method signatures. However, MOTIF-C++ requires the manual specification of the class to instantiate to deal with parametric and subtype polymorphism. For example, if a method declares parameters of an abstract type (e.g., isAutonomous receives one parameter of the abstract type Mission), the end-user needs to specify the additional parameters required to instantiate the concrete class to be used to exercise the method under test (Line 9).
Note that implementing such additional instructions is much less labor intensive than manually identifying the inputs that kill the mutant. Indeed, the proper initialization of the objects required for testing simply consists of the invocation of constructor methods. In contrast, the identification of the inputs that kill a mutant require a comparison between the mutated and the original functions to determine which inputs makes their output differ. Further, we expect that end-users can easily update MOTIF’s test utility classes by copying the code used in unit test cases, except for assignments to primitive variables, which shall rely on the data provided by the fuzzer. Indeed, in unit test cases, like in MOTIF’s test utility class, all the objects required by the method under test shall be instantiated; therefore, the instructions to be manually-specified in the test utility class can be a copy of the ones used in the manually implemented test cases, except for assignments to primitive variables.
Last, we consider the need for the manual modification of test utility classes a limitation of the MOTIF implementation rather than the approach itself. Indeed, it is possible to integrate an additional lightweight static analysis step into MOTIF to automate the generation of test drivers. Indeed, as demonstrated by approaches for the generation of fuzz drivers for API fuzzing FuzzGen , it is generally feasible to implement a lightweight static analysis parser that automatically generates sequences of function calls by copying the content of existing programs (test cases, in our context); however, implementing such a parser goes beyond the objective of this paper which, for the C++ part, is concerned with proposing and assessing a solution to determine when a mutated method is killed (i.e., through a test utility class and a copy of the mutated class) rather than delivering a production-ready tool.
Concluding, although MOTIF-C++ delegates the handling of several C++ features to the manual tuning of the automatically generated test utility class, the automated generation of both the fuzzing driver and the test utility class still significantly decreases manual effort when compared to manual mutation testing.
5 Empirical Evaluation
We address the following research questions:
RQ1. What fuzzer configuration leads to best results in MOTIF? Fuzzing results can vary depending on the configurations applied to the fuzzers, which include the selection of compilers, sanitizers, and coverage metrics. Executables’ speed depend on compilers’ code optimization capability and affect fuzzing effectiveness; indeed, quicker executions lead to more inputs being tested and, likely, to more mutants being killed. Sanitizers are helpful in detecting invalid behaviors, such as memory overflow and arithmetic overflow during software execution; consequently, they may facilitate mutants detection (e.g., mutants leading to such invalid behaviours) although their overhead may reduce the number of generated test inputs and, in turn, mutation testing effectiveness. Coverage metrics are used to determine when an input should be kept in the fuzzing queue, and can affect input generation and mutants killing; for example, specific metrics may facilitate the discovery of unexplored paths, which may include the mutated instructions. Through this research question, we aim to determine what configurations would be optimal for MOTIF.
RQ2. How does mutation testing based on fuzzing compare to mutation testing based on symbolic execution, for software where the latter is applicable? Since certain CPS units may still satisfy the assumptions of SE approaches (e.g., absence of floating-point instructions and black-box components), we aim to assess what approach performs better in such cases.
RQ3. How does MOTIF’s seeding strategy contribute to its results? MOTIF kills mutants either through the generated seeds or through fuzzed inputs; we therefore aim to assess how the two strategies individually contribute to MOTIF results in order to determine if fuzzing is indeed useful.
RQ4. Does hybrid-fuzzing improve the effectiveness of MOTIF? We aim to assess hybrid-fuzzing approaches because they demonstrated to be effective in increasing the code coverage obtained through fuzzing and thus may improve mutants detection. Fuzzing can be an advantage to exercise conditions that are easy to satisfy (e.g., branch conditions controlled by a relational operator in a basic clause, such as x>0) whereas symbolic execution helps satisfy narrow branch conditions (e.g., joining multiple clauses, such as x>0 && x<3yun2018qsym .
RQ5. Is it cost-effective to reuse mutant-killing inputs? Since mutants are often redundant 333 A mutant is considered redundant if it can be killed by the same input that kills another mutant, even though they may differ syntactically shin2017theoretical ., multiple mutants can be killed by the same inputs. Consequently, inputs generated by MOTIF that successfully kill some mutated versions of a function might be used to test other mutated versions that remained live. However, the execution of additional inputs increases mutation testing time. This research question aims to assess the tradeoff between mutation score improvement and increased testing time due to reusing mutant-killing inputs.
5.1 Subjects of the study
Table 3: Subject artifacts.
Subject Open-source LOC # Test cases Statement coverage Mutation score (MS)
MLFS Yes 5,402 4,042 100.00% 81.80%
LIBU No 10,576 201 83.20% 71.20%
ASN1lib Yes 7,260 139 95.80% 58.31%
ESAIL No 2,235 384 95.36% 65.36%
S5 No 54,696 36 62.23% 64.13%
To address our research questions, we considered software deployed on space CPS (satellites) currently in orbit. This included (a) MLFS, the Mathematical Library for Flight Software MLFS , which complies with the ECSS criticality category B ecss40C ; ecss80C , (b) LIBU, which is a utility library developed by GomSpace and used in NanoSatellites, (c) ASN1lib, a serialization/deserialization library, (d) ESAIL, a subset of the control software of a micro-satellite developed by LuxSpace to track ships worldwide, and (e) Sentinel-5 UVNS L1b Prototype Processor (S5, for brevity), which is ground software developed by Huld to process radiation data received from the Sentinel 5 satellite instruments. ASN1lib has been generated with ASN1SCC from a test grammar provided by ESA. ASN1SCC is a compiler that generates C/C++ code suitable for low resource environments ASN1CC ; ASN1:paper . All the subjects are implemented in C except S5, which is implemented in C++.
Our software subjects are provided with test suites whose code coverage is reported in Table 3. Most test suites do not achieve 100% statement coverage because they include components that are tested with specific hardware not available to us; therefore, we generated mutants only for the covered statements. We generated mutants with MASS Oscar:TSE ; MASSTOOL ; specifically, we rely on all the mutation operators supported by MASS (i.e., the sufficient set offutt1996experimental and the deletion set delamaro2014designing ; andrews2005mutation ), which proved effective in previous experiments on similar subjects. We excluded mutants that are identified as equivalent or duplicate according to trivial compiler equivalence methods Oscar:TSE . The last column in Table 3 provides the mutation score (MS) for our case study subjects; it corresponds to the proportion of mutants detected by the test suite. The highest mutation score is observed with MLFS, whose test suite achieves MC/DC adequacy chilenski1994applicability . The lowest mutation score is observed with ASN1lib, which is automatically generated by ASN1SCC using a grammar-based approach ASN1:paper . Our subjects’ mutation scores are in line with empirical investigations reporting mutation scores ranging from 55% to 95% Ramler2017 ; delgado2018evaluation , for CPS software.
To perform test data generation, we rely on the mutants not killed by the original test suites. We assume that the live mutants are not equivalent (i.e., produce the same outputs for every input) to the original software. Although this could be an under-approximation, it does not introduce bias in the comparison between the different approaches considered (i.e., SEMuP and MOTIF configured with different fuzzers and options) because all of them cannot kill equivalent mutants. Further, two mutants $m_{a}$ and $m_{b}$ can also be duplicates (i.e., they lead to the same outputs for every input) or subsumed (i.e., $m_{a}$ is killed by a superset of the test cases killing $m_{b}$). However, the identification of test inputs that kill mutants is a precondition to determine if mutants are duplicate or subsumed Shin:TSE:DCriterion:2018 ; for this reason, including duplicate and subsumed mutants should not introduce bias in the comparison of the approaches. In other words, a mutation testing approach should easily kill mutants that are either duplicates or subsume other killed mutants; if it does not happen, it is correct to penalize such an approach in the empirical evaluation.
The number of live mutants for each subject is: 443 for LIBU, 1,347 for ASN1lib, 3,891 for MLFS, 581 for ESAIL, and 99 for S5. For LIBU, MLFS, and ESAIL, we configured MOTIF to generate arrays of a specific size and, for void pointers (i.e., void *), to create variables or arrays considering the expected data type. Regarding S5, the relatively limited number of mutants considered is due to the need for the manual editing of most of the fuzz drivers generated by MOTIF (see Section 4). Indeed, advanced C++ features, including abstract classes, templates, friend functions, and unnamed namespaces, as well as the standard library implementations, render the drivers automatically generated by MOTIF ineffective. For example, vector<>, one of the standard template libraries (STL), uses pointers to trace the stored data, which renders the memcopy-based solution adopted by MOTIF to instantiate objects inappropriate (pointer addresses cannot be filled with data generated by the fuzzer). To address such limitation, we manually modified the fuzzing drivers automatically generated by MOTIF. Precisely, in total, to kill 99 mutants, MOTIF generated 25 drivers, 21 of which had to be manually modified. However, 14 drivers required a simple modification such as instantiating a template class (e.g., vector<int>), while seven drivers required an initialization step for the class under test, which we copied from the unit test cases provided by S5 developers. In addition, recall that, as discussed in Section 4.1, the need for such manual modifications is a limitation of the MOTIF implementation, not the approach, and can be overcome through an improved static analysis step. Further, compared to manual mutation testing, the automated generation of both the fuzzing driver and the test utility class still significantly decreases manual effort.
5.2 Experimental setup
We performed our experiments using a prototype implementation of MOTIF MOTIF ; MOTIFTool .
To address all our RQs, as fuzzer for MOTIF, we selected AFL++ because it is the fuzzer that performed better in terms of code coverage, according to recent benchmarks in the literature FuzzBenchPaper ; FuzzbenchReport ; GreenFuzz ; moreover, along with HonggFuzz honggfuzz , it is the fuzzer that maximizes fault coverage in another recent benchmark ICST:22:Sarro .
To address RQ2, and compare MOTIF with a mutation testing approach relying on symbolic execution, we modified the MOTIF pipeline to enable test generation with SEMu; we call such pipeline SEMuP. At a high level, SEMuP follows the same steps of MOTIF, with differences concerning how input and output variables are defined.
For SEMuP, in Step 1, we generate SEMu drivers enabling symbolic execution. An example SEMu driver generated for function T_POS_IsConstraintValid is shown in Listing 9. These drivers must specify what are the input parameters to be treated symbolically (see Line 8 in Listing 9), a task performed by the end-user. SEMu drivers do not include explicit comparisons between the outputs of the mutated and the original function because such comparison is handled by SEMu when symbolically executing the original and the mutated functions in parallel (see Section 2.3.1). Precisely, the SEMu driver invokes only the function under test and prints to standard output the data values that should be considered to determine if a mutant has been killed. Similar to MOTIF, SEMu also requires end-users to manually specify how to process data values belonging to data structures referenced with pointers. SEMuP does not include a Step for the generation of seed input (i.e., MOTIF’s Step 2). It includes a step (corresponding to MOTIF’s Step 3) to compile the mutated function and the SEMu drivers with LLVM, followed by a step, corresponding to MOTIF’s Step 4, for the execution of SEMu and the processing of its logs to determine killed mutants.
1int main(int argc, char** argv){
2 // Declare variable to hold function returned value
3 _Bool result;
4 // Declare arguments and make input ones symbolic
5 T_POS pVal;
6 int pErrCode;
7
8 klee_make_symbolic(&pVal, sizeof(pVal), "pVal");
9 // Call function under test
10 result = T_POS_IsConstraintValid(&pVal, &pErrCode);
11 // Print output data
12 printf("pErrCode = %
13 printf("result = %
14 return (int)result;
15}
Listing 9: Example SEMu driver corresponding to the fuzzing driver in Listing 1.
To address RQ3, to perform hybrid fuzzing, we selected SymCC SymCC , whose performance is similar to that of the more recent tools Fuzzolic, SymSAN, and SymQEMU SYMSAN . We excluded SymSAN and its recent extension Marco Marco:ICSE:2024 because SymSAN requires 64-bit compilation, which is not feasible with some of our subjects. We excluded Fuzzolic and SymQEMU since they rely on a version of QEMU that cannot run on some of our subjects (e.g., ESAIL). Therefore, in our response to RQ3, the benefits of hybrid fuzzing can be considered to be a lower bound.
To leverage hybrid fuzzing, we extended MOTIF’s Step 3 to also compile the fuzzing driver and SUT with SymCC’s compiler, in addition to AFL++’s. In Step 4, MOTIF executes the two compiled executable drivers together, following AFL++ procedures for hybrid fuzzing with SymCC [^4] Guideline for SymCC: https://github.com/AFLplusplus/AFLplusplus/tree/ stable/custom_mutators/symcc. When AFL++ discovers new interesting inputs according to its coverage metrics, SymCC is triggered to perform symbolic execution. The inputs found by SymCC are then merged into the queue of AFL++. Once either tool finds inputs killing the mutant, fuzzing stops and MOTIF proceeds with Step 4’s post-processing activity.
In our experiments, we test mutants in parallel, by leveraging the multiple nodes of an HPC infrastructure HPC . This decision depends on the observation that leveraging parallelism (e.g., by relying on Cloud solutions) is a cost-effective choice in the case of complex CPS with many live mutants. Parallel testing of mutants is enabled by the fact that, in contrast to from MU2 MU2 , MOTIF does not test live mutants with the same inputs. Therefore, MOTIF can simply be deployed on different nodes and executed. However, we assess (RQ4) how MOTIF’s effectiveness can be improved by reusing the mutant-killing inputs obtained in the different, parallel executions.
To account for randomness, we executed each approach (i.e., different MOTIF setups and SEMuP) ten times for each subject. For each mutant, we executed each approach for 10,000 seconds, which we determined, in a preliminary study, to be sufficient for SEMuP to maximize the percentage of killed mutants. Precisely, for fuzzing and hybrid fuzzing, we allocate 10,000 seconds of time budget and stop fuzzing once this is reached. But for SEMuP, we allocate 10,000 seconds to the symbolic execution process, which means that, after the timeout, if the mutant has not been killed yet, SEMuP still tries to generate test inputs using the path conditions traversed so far, which leads to an execution time for SEMuP that is slightly higher than others (around 650 seconds more).
5.3 RQ1 - Fuzzer configurations
5.3.1 Design
Since we measure effectiveness in terms of percentage of killed mutants, to address RQ1, we compare the percentage of mutants killed when different fuzzer configurations are selected; specifically, we compare the configuration parameters belonging to three distinct dimensions: compilers, sanitizers, and coverage metrics. For each configuration considered in our experiment, we execute MOTIF for at most 10,000 seconds for each mutant and track the time required to kill each mutant. We compare distinct approaches in terms of number of mutants killed (i.e., mutation score, MS) over time and evaluate the significance of the difference by relying on the Mann–Whitney U-test, a non-parametric test.
The first comparison focuses on the impact of the compiler choice on MOTIF’s effectiveness. We consider two compilers: GCC and Clang. GCC is the default compiler for most software projects in several domains (e.g., space). Clang is a recent compiler that is gaining popularity across industries and is particularly appealing for fuzzing because it integrates several optimization techniques leading to faster programs [^5] AFL++ with LLVM: https://github.com/AFLplusplus/AFLplusplus/ blob/stable/instrumentation/README.llvm.md [^6] Clang-Features and Goals: https://clang.llvm.org/features.html. We select Clang version 14, since it supports all sanitizers and coverage options in AFL++, and GCC version 11.
The second comparison is among sanitizers. We consider four sanitizers supported by AFL++: ASAN Konstantin2012 which detects array out-of-bound and invalid memory access, UBSAN UBSAN:Clang19 which detects undefined behavior such as arithmetic overflow, LSAN LSAN:Clang19 which focuses on heap memory leaks, and MSAN Stepanov2015 which focuses on uninitialized memory reads. We do not consider the two other sanitizers supported by AFL++, Control Flow Integrity SANitizer (CFISAN) and Thread SANitizer (TSAN), because they focus on problems specific for object-oriented and concurrent programs that are unlikely caused by code mutations (indeed, the mutation operators considered for our experiments don’t aim at introducing multi-threading faults nor breaking control flows by altering pointer to functions). To perform our experiment, we compile our subjects (fuzzing drivers and SUT) multiple times, one for each sanitizer, with Clang (the best compiler based on our results, as discussed below), and execute MOTIF on each version. We compare the observed results (MS) with the ones obtained with a baseline consisting of the use of Clang without any sanitization options.
In the third comparison, we considered three coverage metric options supported by AFL++, which is motivated by the fact that, since fuzzers are driven by code coverage, coverage metrics affect the fitness of inputs. They are: (1) LAF LAF:Blog , which splits complex comparison operations into multiple conditions [^7] LAF splits conditions with non-strict relational operators such as if (a <= b)doA or if (a >= b)doB) into compound conditions such as if (a == b){ doA } else if (a > b){ doA }). It also splits invocations of strcmp, to compare strings, into an appropriate set of nested if conditions. Last it splits switch blocks into nested if conditions performing comparisons by-by-byte., thus preserving inputs that exercise boundary cases differently than previous inputs; (2) NGRAM Wang2019 which tracks the coverage of sequences of $N$ edges, thus preserving inputs that exercise different sub-paths in a program, and context-sensitive coverage (CTXWang2019 which combines edge coverage with information on call points, thus enabling the fuzzer to distinguish between the execution of the same code from different calling contexts. For LAF, we enabled the splitting of all the types of compound expressions supported by LAF, and for NGRAM, we set the parameter N to 2. To compile our subjects, we rely on Clang.
Table 4: Results obtained by MOTIF with different fuzzer configurations. Each cell reports the average percentage of mutants killed over ten runs by each configuration after 10,000 seconds and $p$-values based on U-test against the baseline.
Experiment group Options LIBU ASN1lib MLFS ESAIL S5
Compilers (baseline: GCC) GCC $48.31\%$ $85.44\%$ $35.52\%$ $38.38\%$ $83.03\%$
Clang $49.68\%$ $(0.0009)$ $86.57\%$ $(0.0002)$ $38.43\%$ $(0.0002)$ $39.04\%$ $(0.0001)$ $82.73\%$ $(0.2641)$
Sanitizers (baseline: Clang) ASAN $49.22\%$ $(0.1081)$ $86.82\%$ $(0.0200)$ $38.29\%$ $(0.0019)$ $39.36\%$ $(0.0048)$ $81.52\%$ $(0.0007)$
UBSAN $51.35\%$ $(0.0003)$ $84.31\%$ $(0.0002)$ $23.78\%$ $(0.0002)$ $38.67\%$ $(0.0001)$ $82.42\%$ $(0.2781)$
LSAN $47.39\%$ $(0.0002)$ $85.78\%$ $(0.0003)$ $37.60\%$ $(0.0002)$ $38.62\%$ $(0.0002)$ $80.81\%$ $(0.0000)$
MSAN $45.56\%$ $(0.0002)$ $83.39\%$ $(0.0002)$ $34.33\%$ $(0.0002)$ $38.30\%$ $(0.0001)$ $6.06\%$ $(0.0000)$
Coverage metrics (baseline: Clang) LAF $49.79\%$ $(0.6158)$ $88.84\%$ $(0.0002)$ $40.09\%$ $(0.0002)$ $38.73\%$ $(0.0004)$ $82.53\%$ $(0.3006)$
NGRAM $50.85\%$ $(0.0016)$ $86.27\%$ $(0.0023)$ $38.27\%$ $(0.0045)$ $38.97\%$ $(0.3446)$ $83.64\%$ $(0.0005)$
CTX $50.89\%$ $(0.0030)$ $86.49\%$ $(0.3989)$ $38.29\%$ $(0.0070)$ $38.98\%$ $(0.5518)$ $83.74\%$ $(0.0001)$
n.nnand n.nn%: significantly outperformed the baseline with $p\leq 0.01$ and $p\leq 0.05$, respectively
n.nnand n.nn%: significantly underperformed, compared to the baseline, with $p\leq 0.01$ and $p\leq 0.05$, respectively
5.3.2 Results
Table 4 reports MOTIF’s performance, for each subject, when different fuzzer configurations are selected. Each row shows the average percentage of live mutants killed by MOTIF over ten runs after 10,000 seconds, for a given fuzzer configuration. $p$-values are shown in brackets, capturing the statistical significance of the difference between the selected configuration and the Clang baseline; for the compiler selection assessment, $p$-values simply capture the significance of the difference between the results obtained using the two available options. A green background highlights cases where the selected configuration significantly outperformed the baseline; a red background highlights the opposite. Cases not leading to any significant difference have a white background.
The compilers’ comparison results show that Clang outperforms GCC by 1.37 percentage points (pp) in LIBU, 1.13 pp in ASN1lib, 2.91 pp in MLFS, and 0.65 pp in ESAIL. For each subject of these four subjects, the results obtained with the two compilers are significantly different at every minute. The better performance of Clang is likely due to the fact that the number of inputs generated by AFL++ with Clang is 4 to 20 times higher than the one observed with GCC. Indeed, search algorithms (i.e., AFL++) are more likely to explore a larger portion of the input space when they generate more inputs. The larger number of inputs generated with Clang can be explained by Clang generating executable programs (in our case, fuzz drivers) that are quicker to execute than GCC ones and therefore AFL++ can try more inputs with Clang than with GCC, for the same test budget. In contrast to the above, for S5, Clang performed worse than GCC by 0.30 pp but the difference is not statistically significant, enabling us to conclude that overall Clang is the best choice for MOTIF.
The sanitizer results in Table 4 show that MOTIF equipped with sanitizers tends to perform worse than the baseline, MOTIF without any sanitizer. ASAN led to the highest number of cases with improvements; indeed, it slightly outperformed the baseline by 0.25 pp ($p<0.05$) in ASN1lib and by 0.33 pp ($p<0.01$) in ESAIL. UBSAN, instead, led to the largest improvement, with a proportion of killed mutants higher than the baseline by 1.67 pp ($p<0.01$) in LIBU; however, it performs worse than the baseline for most of the subjects.
The primary reason why MOTIF with sanitizers cannot outperform the baseline is the overhead of the instrumented code introduced by sanitizers to detect invalid behaviors. Such overhead reduces the number of inputs generated by AFL++, consequently decreasing the number of mutants killed, even though sanitizers may kill some mutants not killed without them. In our investigation, the worst performance is obtained with MSAN, which led to 96.77% fewer inputs than the baseline, on average, and showed significantly lower performance. The worst MSAN result is obtained with S5, likely because of poor support for C++ libraries (e.g., MSAN led to crashes whenever it encounters a standard library function such as std::map and std::set). ASAN, instead, led to 61.5% fewer inputs, on average, but enabled the detection of a large number of additional mutants thus compensating the loss due to execution cost; consequently, ASAN achieved results that are comparable to those obtained without sanitizer. ASAN effectiveness in killing mutants mainly depends on its capability to detect out-of-bound memory accesses caused by mutants; such mutants are harder to detect otherwise because the sufficiency condition (see Section 2.3) is hard to meet without a sanitizer (the effect of the out-of-bound access shall affect the values assigned to an output variable). Further, ASAN prevents the generation of test inputs that kill mutants but are not valid; those are inputs that violate function preconditions and cause out-of-bound accesses in the original function under test, and are thus discarded by MOTIF. Different from ASAN, UBSAN prevents the testing of several mutants because it detects invalid behaviors in the original function (on average, 653.8 fuzzing drivers failed when executing the original function), which need to be fixed before proceeding with mutation testing. A representative example is provided by MLFS, where UBSAN detects several arithmetic overflows caused by shift operators applied to variables that may have negative values. Although UBSAN may have detected potential bugs in the SUT, by terminating the execution of the original function, UBSAN prevents mutation testing. In summary, ASAN is the only sanitizer that is beneficial to MOTIF for two of the subjects (i.e., prevents false positives without largely affecting effectiveness).
Coverage metrics results in Table 4 show that their effectiveness vary across subjects. Specifically, MOTIF with LAF outperformed the baseline (i.e., MOTIF with AFL++ using the default coverage options) by 2.27 pp and 1.66 pp, with $p$-value $\leq 0.01$, in ASN1lib and MLFS, respectively. With these two subjects, LAF likely improves reachability (i.e., generating inputs that reach the mutated statement) because they present several conditional statements that can be split into simpler conditions; indeed, ASN1lib contains compound conditions that check input validity, while MLFS, relies on several signed and floating point comparisons. In ESAIL, LAF performed significantly worse than the baseline by 0.31 pp; however, the practical impact of such decrease in performance is very limited (2 fewer mutants killed, in the worst case).
MOTIF with NGRAM and CTX outperformed the baseline in both LIBU and S5. In LIBU, we observe an improvement of 1.17 pp and 1.21 pp for NGRAM and CTX, respectively ($p$-value $\leq 0.01$). In S5, the improvements are 0.91 pp and 1.01 pp, respectively ($p$-value $\leq 0.01$). In LIBU, the improvements are likely due to LIBU including longer call sequences whose results depend on state variables. In S5, the improved performance is likely due to C++ language features (i.e., overriding and polymorphism), resulting in method invocations that behave differently depending on the call context. In the worst case, NGRAM and CTX, decrease MOTIF’s performance by 0.30 pp (NGRAM on ASN1lib) and 0.14 pp (CTX on MLFS), with an average of 4 to 6 fewer mutants being killed, a small impact in practice. To summarize, it is not possible to identify a coverage metric that works best with all the subjects; however, since the differences are limited, engineers may select the option that improves testing results in critical code units (e.g., LAF because it improves MS in mathematical functions).
Although our results did not help identifying a sanitizer and coverage metric configuration that provides effectiveness improvements across all subjects, for the sake of simplifying the remaining experiments, we identify a configuration for MOTIF that is likely to provide improvements in most subjects. It consists of combining Clang, ASAN, and LAF and we call such configuration Best, . We selected ASAN because it performs similarly to the baseline in terms of percentage of killed mutants but also prevents the generation of inputs that are invalid and cause memory access errors. We selected LAF because it leads to the largest number of killed mutants, across all subjects (3283 versus 3186 for NGRAM and 3190 for CTX), and further leads to the largest effectiveness improvement (+2.27 pp), and also improves the subject with the most effective test suite (i.e., MLFS, which achieves MC/DC and 81.80% MS).
(a) LIBU
(b) ASN1lib
(c) MLFS
(d) ESAIL
(e) S5
Figure 3: Comparison between MOTIF with GCC and MOTIF with the Best fuzzer options.
To ensure that the selected configuration options do not interfere, we performed ten additional executions of MOTIF configured with the Best settings, with each subject, and compare the observed results with MOTIF relying on the default AFL++ options (i.e., relying on GCC without additional sanitizers or coverage metrics). Figure 3 shows the percentage of mutants killed by MOTIF after each second, when using GCC and Best, for each subject. Each line captures the average percentage observed in ten runs, with the shaded area capturing the upper and lower bounds across those runs. The vertical dashed line shows the time budget for the experiment. With Best, the percentages of killed mutants are 49.84% for LIBU, 88.74% for ASN1lib, 39.85% for MLFS, 39.50% for ESAIL, and 82.53% for S5. These performance results are significantly better than those of our baseline (i.e., MOTIF with GCC) for four subjects: it increases the percentage of killed mutants by 1.53 pp in LIBU, 3.30 pp in ASN1lib, 4.34 pp in MLFS, and 1.12 pp in ESAIL. For S5, the performance is similar to that of the baseline. Further, we manually inspected the mutants killed only by the Best configuration and observed that 70% of the generated test cases enable exercising boundary cases missed by the original test suite, which include access to first/last items of arrays and use of boundary values in conditional statement. We can thus conclude that Best helps improve the quality of test suites because in safety-critical systems it is desirable to exercise such situations. We can thus conclude that combining Clang, ASAN, and LAF leads to the best results in MOTIF.
We leave the identification of solutions to overcome the mutation score plateau (see Figure 3) to future work; however, it could also be partially due to the presence of equivalent mutants.
5.4 RQ2 - Fuzzing vs Symbolic Execution
5.4.1 Design
We compare fuzzing and symbolic execution in terms of cost-effectiveness. The effectiveness of an automated mutation testing tool can be measured in terms of the proportion of live mutants killed. Its cost is determined by the time required to kill the mutants; indeed, lengthy test data generation may delay the testing process and increase the usage of computing resources. Though cost is also driven by the time required to manually inspect test outputs, MOTIF and SEMuP should require the same manual inspection time because they invoke the same functions under test and print out the same output values. Therefore, regarding cost, we focus on execution time and thus compare cost-effectiveness in terms of live mutants killed for different time budgets.
We could not consider MLFS to address RQ2 because it works mainly with floating point arguments, which are not supported by KLEE. An old version of KLEE addresses floating point variables but it is not integrated into SEMu. We could not consider ESAIL and S5 because of the lack of KLEE support for multi-threading libraries and C++. We therefore focus on LIBU and ASN1lib; however, for LIBU we considered only four out of 27 source files, because all the other source files included I/O operations, which are not supported by KLEE/SEMu, or cannot be compiled into LLVM bitcode. This leads to 1,347 live mutants for ASN1lib and 153 for LIBU. For MOTIF, we considered the Best configuration identified in RQ1.
5.4.2 Results
(a) LIBU
(b) ASN1lib
Figure 4: Percentage of mutants killed by MOTIF and SEMuP.
Figure 4 depicts the percentage of live mutants killed by MOTIF and SEMuP for LIBU (4(a)) and ASN1lib (4(b)), respectively. Each line represents the average percentage from ten runs, with the shaded area around each line indicating the upper and lower bounds across those runs. The vertical dashed line means the time budget for each experiment. At that point, SEMuP stops exploring paths and generates inputs that satisfy the current path condition, which, sometimes, is sufficient to identify inputs that kill mutants. We observe a rapid increase in the number of mutants killed by SEMuP for the ASN1lib, which includes paths with several nested conditions.
The plots show that MOTIF outperforms SEMuP. After 10,000 seconds, MOTIF kills between 117 (76.47%) and 120 (78.43%) mutants for LIBU (avg. is 118.20, 77.25%) and between 1,190 (88.34%) and 1,200 (89.09%) for ASN1lib (avg. is 1,195.30, 88.74%). In contrast, SEMuP kills 41 (26.80%) to 42 (27.45%) mutants for LIBU (avg. is 41.2, 26.93%) and 1,017 (75.50%) to 1,018 (75.58%) for ASN1lib (avg. is 1,017.8, 75.56%). On average, across the ten runs, MOTIF kills a percentage of mutants that is 50.33 percentage points (pp) and 13.18 pp higher than SEMuP’s, for LIBU and ASN1lib, respectively.
The difference between MOTIF and SEMuP is significant at every timestamp, based on a U-test ($p<0.01$). For example, after one minute, MOTIF kills, on average, 110.5 (LIBU) and 1,045.00 (ASN1lib) mutants, while SEMuP kills 29 (LIBU) and 924.6 (ASN1lib) mutants. For both LIBU and ASN1lib, MOTIF quickly reaches a near plateau; in LIBU, MOTIF reaches the plateau quicker because of LIBU’s simple control logic.
Though MOTIF outperforms SEMuP, they show some degree of complementarity, which justifies the integration of hybrid fuzzers in MOTIF (see Section 2.2). If we consider the best run of each approach, in the case of ASN1lib, MOTIF kills 257 (19.08%) mutants not killed by SEMuP, while SEMuP kills 75 (5.57%) mutants not killed by MOTIF. In the case of LIBU, MOTIF kills an additional 78 (50.98%) mutants on top of the mutants killed by SEMuP. We manually inspected some of the mutants and noticed that SEMuP is sometimes better at generating inputs that satisfy narrow, simple constraints. However, such a characteristic is more useful for ASN1lib, which mainly performs boundary checks for nested data structures, rather than the utility library. On the other hand, MOTIF is better when SEMuP fails to solve complex constraints. For example, for LIBU, SEMuP could not kill 52 mutants affecting a conditional statement with 24 bitwise operations, 44 mutants affecting a conditional statement with 13 conditions expressed using inequalities, and 5 mutants affecting the size of the buffer used in snprintf statements. Finally, MOTIF enabled the discovery of four bugs that were confirmed by developers: two concern missing checks for out-of-domain numerical inputs, the other two are an integer overflow affecting operations on a numerical data structure with multiple fields. We detected the missing checks by observing that the generated test cases include an out-of-domain input but do not result in an execution error flag being set (likely they were not detected by test cases because the valid domain is underspecified in specifications). The integer overflow was discovered by observing that the result of a sum of two large numeric items led to a lower number. The integer underflow can be noticed by observing that the difference between a zero-filled data instance and a data instance with small positive numbers result in large positive numbers. SEMuP discovered only three of them.
5.5 RQ3 - Seeding effectiveness
5.5.1 Design
To discuss how MOTIF’s seeds contribute to mutation testing results, we focus on the proportion of mutants killed with seed inputs in the experiments performed to address RQ1 with the MOTIF’s Best configuration.
5.5.2 Results
MOTIF’s seed inputs contributed to killing mutants as follows: 22.2 (5.08%) for LIBU, 304 (22.57%) for ASN1lib, 81 (2.08%) for MLFS, 69 (11.88%) for ESAIL, and 39 (39.39%) for S5. However, although in all the subjects except LIBU seed inputs killed the same mutants across the ten runs, in LIBU the number of killed mutants varied between 22 and 23 because of a non-deterministic function that computes the time difference before and after invoking the sleep() function.
The percentage of mutants killed by seed inputs largely depends on the nature of the functions under test. For mutants in MLFS and LIBU, such percentages are low because they mainly alter mathematical operations whose mutants are killed with inputs satisfying complex constraints. For ASN1lib, ESAIL, and S5, the proportion of mutants killed is higher because the mutants modified lines that affect the output observed with any input value. Please note that seed inputs do not introduce bias in RQ2 results since SEMuP kills most of the mutants killed by seed inputs ($267/304$ for ASN1lib and $/1$ 1 for LIBU).
Concluding, although the selected seed inputs help kill mutants, the contribution of the fuzzing process is significant with, at the very minimum (S5), 52.27% (i.e., $100\%-\frac{39.39\%/}{82.53\%}$) of the killed mutants being killed by fuzzing.
5.6 RQ4 - Applying Hybrid-fuzzing
5.6.1 Design
We compare MOTIF with MOTIF-Hybrid, our MOTIF extension integrating AFL++ with SymCC to leverage hybrid fuzzing for mutation testing (see Section 5.2). To ensure a fair comparison, we applied the Best fuzzing configuration identified in RQ1 to both approaches. As for RQ2, we compare the two approaches in terms of cost-effectiveness, by reporting on the MS and the time taken to kill mutants.
We considered four subjects: LIBU, ASN1lib, MLFS, and ESAIL. We excluded S5 because SymCC cannot successfully compile C++ code relying on the STL library. Also, since SymCC fails to correctly compile code chunks with increment and decrement operators for boolean, we excluded two ESAIL mutants. Finally, we executed the two approaches ten times on each subject for 10,000 seconds.
(a) Mutation score
(b) Kill time (seconds)
Figure 5: Results observed when applying MOTIF with AFL++ (MOTIF) and MOTIF with SymCC (MOTIF-Hybrid).
5.6.2 Results
Figure 5 presents boxplots providing the distributions of MS (5(a)) after 10,000 seconds and the time taken to kill mutants (5(b)), across the ten experimental runs.
MOTIF-Hybrid increases the percentage of killed mutants in ASN1lib by up to 1.37 pp (18.5 mutants) on average, while the performance differences for the other subjects are not statistically significant.
As for the time taken to kill mutants, we report that the mutants in ASN1lib and MLFS were killed significantly faster with MOTIF-Hybrid. The median per mutant for ASN1lib and for MLFS is 142.00 vs. 208.66 seconds and 214.60 vs. 253.08 seconds, respectively, when compared with MOTIF. In LIBU and ESAIL, the difference is not significant, however, with the median observed for MOTIF-Hybrid with LIBU being lower (57.13 seconds with MOTIF-Hybrid vs. 66.13 seconds with MOTIF).
Table 5: Complementarity between MOTIF-Hybrid, MOTIF, and SEMuP in ASN1lib: number of mutants killed and not killed by the selected approaches, for their best execution.
MOTIF-Hybrid SEMuP
KILLED LIVE KILLED LIVE
SEMuP KILLED 955 63 - -
LIVE 267 62 - -
MOTIF KILLED 1188 12 943 257
LIVE 34 113 75 72
Additionally, we manually investigated the complementarity between MOTIF-Hybrid, MOTIF, and SEMuP in ASN1lib, which, in RQ2, showed that MOTIF and SEMuP are complementary. Table 5 shows the number of killed and live mutants observed in the best run of each approach. MOTIF-Hybrid killed 267 mutants (19.82%) that were not killed by SEMuP. However, MOTIF-Hybrid killed only 12 of the 75 mutants killed by SEMuP but not by MOTIF; the fact that SymCC cannot kill all the mutants killed by SEMuP shows that SymCC presents some limitations compared to SEMuP (e.g., it may need to augment the set of inputs executed symbolically). However, if we subtract the mutants missed by MOTIF-Hybrid from the additional mutants killed by MOTIF-Hybrid compared to SEMuP and MOTIF, we observe that MOTIF-Hybrid still kills 204 more mutants than SEMuP and 22 more mutants than MOTIF, thus remaining the most effective choice.
Concluding, MOTIF-Hybrid is the best approach for mutation testing since it is faster at killing mutants and has a higher MS in half of the subjects, while getting an equivalent MS with the others, when technical limitations do not prevent its adoption (e.g., in C++ projects like S5). MOTIF-Hybrid could therefore be a practical solution for large projects resulting in a significant number of mutants. However, the setup required for MOTIF-Hybrid is more complex than the one required for MOTIF; indeed, it is necessary to set up additional configurations for the AFL fuzzer and to compile the fuzzing drivers twice, with the AFL compiler and with SYM-CC. Consequently, MOTIF-Hybrid is less likely to be adopted in practice, in favor of MOTIF.
5.7 RQ5 - Reusing inputs
Table 6: Comparison MOTIF with Best set of fuzzing options and results after reusing inputs for five selected subjects
Subject Target mutants 3 c # of killed mutants (Avg. of 10 runs) 3 c # of inputs for functions (10 runs) 3 c Reuse time (50 runs, seconds) 3 c Kill time (50 runs, seconds)
FUZZED REUSED DIFF Mean Min Max Mean Min Max Mean Min Max
LIBU 325 61.78% 62.71% 0.92 pp (0.0048) 299.49 1 3166 5.72 0.01 264.84 0.21 0.02 5.02
ASN1lib 762 80.09% 83.88% 3.79 pp (0.0002) 143.69 4 779 9.13 0.10 854.79 0.12 0.01 1.06
MLFS 3623 42.50% 43.57% 1.08 pp (0.0002) 125.65 1 1267 12.51 0.01 353.51 0.11 0.01 4.14
ESAIL 418 54.90% 55.02% 0.12 pp (0.2515) 246.78 1 1398 484.40 0.23 835.10 0.24 0.14 1.06
S5 28 88.21% 89.29% 1.07 pp (0.0767) 11.33 7 28 1.76 0.00 6.74 0.10 0.08 0.15
n.nn pp%: outperformed the baseline with $p\leq 0.01$,
5.7.1 Design
RQ5 aims to assess the tradeoff between mutation score improvement and increased testing time observed when reusing inputs killing mutants to kill other mutants not successfully killed by MOTIF. For each mutated function with at least one mutant killed by MOTIF, we rely on the mutant-killing inputs generated by MOTIF to test the mutants not successfully killed by MOTIF. Precisely, we collect the fuzzed files that kill mutants in a function and, after excluding duplicates, provide them as inputs to the fuzzing driver for a live mutant of the same function. We also apply the MOTIF post-processing step to avoid false positives (see Step 4 in Section 3.4).
We considered all the five subjects used for RQ1, and the mutant-killing inputs obtained with the Best configuration. To account for randomness, we apply the approach to the results of all the ten runs considered for RQ1.
To discuss effectiveness, we compare the MS obtained by reusing inputs with the MS obtained by MOTIF after 10,000 seconds.
As for cost, we measure, for each mutant, the time taken to execute fuzzing drivers with all inputs and the time taken until the mutant is killed. Since the order of inputs may affect execution time, we repeat our experiment 50 times, after shuffling each input set.
5.7.2 Results
Table 5.7 presents our results. Column target mutants reports the number of mutants that belong to the functions that have at least one killed mutant. Column FUZZED reports the MS obtained by MOTIF on average over 10 runs. Column REUSED reports the MS obtained after reusing inputs to kill additional mutants. Column DIFF shows the differences between FUZZED and REUSED, and reports on the significance of the difference (we highlight significant improvements). The next three sets of columns report statistics (mean, min, and max) on the number of inputs available for reuse across target functions, the time required to test each mutant with the reused inputs, and the time taken to kill a mutant.
The experiment results indicate that MS can be improved by up to 3.79 pp when reusing inputs, with significant improvements in most of the subjects, which indicates that input reuse is beneficial for MOTIF. The time taken for reusing inputs varies, taking up to 15 minutes (900 seconds) depending on several factors, including the number of inputs, software size, and the type of mutants. For instance, ESAIL has approximately 70% of its mutants within a single function, which leads to a large number of additional executions when reusing inputs. In the case of S5, the larger size causes a longer execution time for fuzzing drivers compared to other subjects, but the maximum execution time is lower due to the number of inputs collected. The execution time is also affected by the type of mutants. Specifically, some mutants took longer than others because they modified stopping conditions in program loops.
Since, in the worst case, input reuse takes 855 seconds, and given that Figure 3 shows that after 9000 seconds MOTIF has already reached the plateau, we can conclude that with a test budget of 10,000 seconds per mutant, it is convenient to rely on input reuse. For lower test budgets input reuse may provide benefits only for a subset of subjects. However, based on our results, up to 5000 seconds of test budget, input reuse is likely to be beneficial; indeed, in all our subjects, after 4150 seconds, MOTIF has killed at most 1 pp mutants less than at 10,000 seconds. Concluding, we suggest executing MOTIF with a budget of 5000 seconds per mutant, and dedicate the last 900 seconds to input reuse.
5.8 Threats to validity
To address threats to internal validity, we manually verified that MOTIF, SEMuP, and MOTIF-Hybrid correctly execute; and further, we manually inspected a large subset of the generated test cases and mutants killed in our experiments. Further, our false positive driver ensures that MOTIF results are not affected by the presence of global variables or, more generally, non-determinism. Although we do not reset global state variables in fuzzing drivers, note that across all experiment runs, out of 32,875 mutants reported as killed by the fuzzing driver, only 125 were false positives (0.38%), thus showing that non-determinism does not undermine the applicability of MOTIF.
Though our results may depend on the specific fuzzer used in our experiments, AFL++ is one of the best performing grey-box fuzzers according to recent benchmarks (see Section 5.2). Further, we assessed MOTIF with a state-of-the art hybrid fuzzing solution. Although alternative hybrid fuzzers were not applicable to some of our subjects (see Section 5.2), they might have led to different and potentially better results; we leave their investigation to future work.
To address generalizability issues, we selected diverse software subjects that are installed and running on space CPS, including satellites currently in orbit: a mathematical library, a utility library, a data serialization component, onboard control software, and ground software. Since they implement a diverse set of features (mathematical operations, serialization, string and time utilities), they strengthen the generalizability of our results. Further, these types of software components are typical in many CPS systems including avionics, robotics, and automotive, thus suggesting the proposed approach may be useful in many sectors other than space.
6 Conclusion
We propose MOTIF, an approach that leverages fuzzing to automatically generate test data for mutation testing of embedded software deployed in cyber-physical systems (CPS). It aims to overcome the limitations of SOTA approaches, which rely on symbolic execution and cannot easily be applied in many contexts, especially CPS ones.
MOTIF is implemented through a pipeline that generates a test driver that processes the input data generated by the fuzzer, provides such input data to the original and mutated versions of a function under test, and determines when the outputs generated by the two functions differ (i.e., the mutant is killed). By monitoring the coverage achieved when executing the original and mutated functions, the fuzzer identifies inputs leading to different behaviors across these functions and, consequently, is driven towards the identification of inputs that kill the mutant.
We performed an empirical evaluation with embedded software deployed on satellites currently in orbit. We empirically determined the fuzzer configurations leading to the best mutation testing results, which consists of relying on the Clang compiler with address sanitization and LAF coverage optimization. In our subjects, such configuration enables killing between 40% and 83% of the mutants. Further, although our seeding strategy contributes to quickly killing mutants, most of the mutants (between 60% and 97%) are killed thanks to the fuzz testing process. We compared MOTIF with a SOTA approach based on symbolic execution, which showed that the percentage of mutants killed by MOTIF is higher than the SOTA approach by 13 and 50 percentage points in our two case studies where symbolic execution is applicable. Our results therefore clearly show that fuzzing should be adopted as the preferred method to use to perform mutation testing. However, we also demonstrated that hybrid-fuzzing, which integrates fuzzing and symbolic execution, leads to slightly increasing the percentage of killed mutants (up to 1.37 pp).
Acknowledgment
The experiments presented in this paper were carried out using the HPC facilities of the University of Luxembourg (see http://hpc.uni.lu). This research was supported by ESA via a GSTP element contract (RFQ/3-17554/21/NL/AS/kkIMPROVE) and by the NSERC Discovery and Canada Research Chair programs. The authors would like to thank Oscar Cornejo and Enrico Viganò for having contributed to the preliminary versions of MOTIF.